# Strategic Tradeoffs Between Humans and AI in Multi-Agent Bargaining

### v3

@@ 6.2.2 Strategic alignment and regret minimization
Figure 4: Example of player chip value trajectories over nine turns of a trading game (3-Chip). To avoid speculation, we do not analyze analyze Player 1’s proposal on Turn 4 (an unaccepted trade), or Player 3’s acceptance on Turn 5 (accepting a negative surplus offer).
Figure 5: Counts of strategic actions across games. Bayesian agents performed more optimal, no-regret actions relative to other populations as game complexity increased. Declining a proposal would not cause forced regret, as this action does not commit any chips. The total number of actions naturally increases with game complexity, reflecting the larger solution space, not necessarily improved decision-making.
Figure 5 shows the classifications of actions across games and populations. [^13] We exclude refined LLMs as they behave similarly to out-of-the-box models; see Appendix F.1. The relative performance of Bayesian agents improves as game complexity increases. Combined with the surplus outcomes in Figure 2, this suggests that Bayesian agents are not necessarily improving, but rather that LLM and human performance degrades as the action space expands.
Bayesian agents generate a higher proportion of optimal (no regret) proposals and fewer regrettable ones compared to LLMs and humans, reflecting more effective surplus-maximizing calculations. However, Bayesian agents also register the most declines—both regrettable and non-regrettable. This indicates that their proposal strategies prioritize their own advantage, without necessarily considering whether those offers will be accepted by others.
In contrast, humans and LLMs exhibit more regrettable proposals and acceptances relative to Bayesian agents. They also exhibit fewer declines overall, indicating that they propose more agreeable offers at the cost of making less optimal actions. Differences in forced regret actions are relatively modest, though LLMs show slightly higher forced regret counts.
7 Discussion
Humans reflect economic and social reasoning.
Humans offered trades consistent with classical fairness norms, proposing a balanced trade ratio. While these proposals were less frequently optimal than that of Bayesian agents, they were also accepted more often. Post-game survey responses consistently emphasized fairness and cooperation as decision drivers, suggesting that humans prioritize social norms, even in one-shot interactions where long-term reputation effects would not affect payoff. [^14] Post-game survey questions and quantitative responses are in Appendix C. However, nuanced human social motivations are difficult to formalize into algorithms, posing a fundamental challenge when delegating coordination and negotiation to artificial agents.
LLMs can perform on-par, but may be concessionary by design.
The latest model, GPT-4o, achieved human-level surplus with minimal prompting, demonstrating flexibility when applied to novel tasks. However, their behavior was compromising and reactionary: they favored value-balanced proposals which were less optimal than Bayesian proposals but more likely to be accepted, and exhibited high rates of forced regret trades, suggesting limited strategic foresight in dynamic scenarios. This conservative play style may arise from several factors: (i) training on cooperative, information-sharing dialogues, where minimizing friction is implicitly rewarded (Wei et al., 2022); (ii) a tendency to generate risk-averse, passive responses compounded by a lack of outcome-driven feedback (Ouyang et al., 2022); and/or (iii) a possible learned aversion to asymmetric outcomes, where generous offers serve as a coordination strategy to secure agreement (Fisher et al., 2011). In adversarial or zero-sum negotiations, this overly concessionary tendency could lead LLMs to systematically overcompromise.
Bayesian agents outperform, but lack social adaptability.
Bayesian agents excelled at surplus-maximizing proposals, aggressively extracting value. This strong performance was constructed through a hand-crafted algorithm specifically aligned to the game’s incentives. In contrast, humans and LLMs relied on their flexible “out of the box” capabilities, lacking such tailored inductive biases.
In practice, Bayesian agents’ extractive, rejection-tolerant strategy might fail in real-world negotiations, especially in repeated interactions requiring trust and reciprocity. Extending Bayesian agents to more socially compatible trading would require richer models of preferences, fairness, and adaptive behavior, which could increase complexity and decrease robustness.
Hybrid models offer promise, but need social reasoning.
We find that humans consider social factors, LLMs favor safe, consensus-seeking strategies, and Bayesian agents excel through narrow, task-specific optimization. These contrasting strengths suggest potential for complementary approaches, where LLMs are augmented with Bayesian tools or planning modules to improve foresight, such as inventory management and lookahead reasoning. However, negotiation is not just a planning problem; success also depends on modeling others’ intentions, social norms, and the unwritten rules of cooperation. Without explicit social reasoning, even hybrid agents will struggle with the social nuances that humans navigate instinctively.
Procedural alignment matters for human-AI interaction.
Furthermore, how agents negotiate is as important as what they achieve. Despite humans and LLMs yielding similar aggregate surplus in some contexts, they exhibited distinct strategies. Crucially, these differences only became visible through a custom experimental design that enabled matched, one-to-one comparisons across populations; simple outcome-level aggregation would have obscured these divergences. As AI agents increasingly participate in human-facing decision-making, focusing solely on efficiency metrics risks missing misalignments in process, intent, and social compatibility.
Efficiency and scale advantages.
A final consideration beyond strategy is scale. LLMs and Bayesian models can resolve negotiations near-instantaneously, unaffected by cognitive load or decision fatigue. LLMs, in particular, are faster to deploy out-of-the-box compared to bespoke Bayesian models; in our case, our out-of-the-box agents required minimal prompt scaffolding around the human instructions, whereas we constructed a custom algorithm for the Bayesian agents. These efficiency gains make automation attractive for high-volume or time-sensitive negotiations, even if strategic or social limitations remain.
Social impacts and risks.
Agents that achieve surplus-yielding outcomes through fundamentally different negotiation styles pose risks for trust, fairness, and social acceptance. If deployed in real-world negotiations, models like LLMs may default to safe, concessionary behavior, leading to suboptimal deals and reinforcing status-quo power dynamics. Conversely, optimization-focused agents like Bayesians may extract disproportionate value, undermining perceptions of fairness and cooperation. These dynamics may lead to unexpected interaction effects in mixed-agent environments. Designing AI agents that are not only strategically capable but also socially compatible will be critical for their long-term acceptance and responsible deployment.
8 Limitations and future work
@@
Our study focuses on a single, stylized negotiation game with static valuations and single shot interactions, conditions that align well with Bayesian agents’ task-specific optimization and may exaggerate their advantage. While humans and LLMs bring general-purpose reasoning to the task, their broader capabilities in dynamic, multi-round, or reputation-driven negotiations remain unexplored. Furthermore, our LLMs used minimal prompting, reflecting realistic deployment but limiting exploration of more strategic behaviors from alternative prompts or fine-tuning.
@@
Beyond performance metrics, our analysis reveals distinct behavioral profiles but does not disentangle whether these arise from reasoning limitations, social norms, or inductive biases inherent to each system. Future work should investigate how humans perceive and trust AI negotiators with different procedural styles. Another promising avenue for future work is to allow agents to communicate with natural language. Finally, developing methods to steer LLM agents towards desired procedural norms explicitly represents a significant ongoing challenge and opportunity.
9 Conclusion
This study demonstrates an empirical approach for comparing agents in dynamic, social settings. By examining human, LLM, and Bayesian behavior in a structured bargaining game, we show how performance measures can conceal fundamentally different strategic approaches. As AI systems are increasingly delegated complex tasks, understanding how their decision-making processes align with human expectations and social values is critical to fostering trust, effective cooperation, and responsible deployment in real-world applications.
@@ Appendix A Algorithms and proofs
A.1 Proof of no dominant strategy
@@
Construct two different scenarios of other agents’ strategies and valuations, denoted by $\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$. These scenarios can differ in 1) how other agents value the goods $\{v_{jg}:j\neq i\}$, and 2) the demands or offers they make (i.e. how $\mathbf{s}_{-i}$ translates into allocations $a_{jg}$).
@@
  • In scenario $B$, suppose the others behave differently (e.g., they no longer place much value on $g^{*}$ but highly value some other good). Now a different specialized strategy $s_{i}^{B}$ might yield a strictly higher payoff (because trading strategy for $g^{*}$ in scenario $A$ is no longer beneficial here).
Contradiction of dominance: Because scenarios $A$ and $B$ are both feasible profiles of opponents’ behavior ($\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$), there is no single strategy $s_{i}^{*}$ that can simultaneously guarantee at least as high a payoff as both $s_{i}^{A}$ and $s_{i}^{B}$ across those distinct opponent profiles. Formally,
$\exists\,\mathbf{s}_{-i}^{A}:u_{i}\bigl(s_{i}^{A},\,\mathbf{s}_{-i}^{A}\bigr)\;>\;u_{i}\bigl(s_{i}^{*},\,\mathbf{s}_{-i}^{A}\bigr),\quad\text{and}\quad\exists\,\mathbf{s}_{-i}^{B}:u_{i}\bigl(s_{i}^{B},\,\mathbf{s}_{-i}^{B}\bigr)\;>\;u_{i}\bigl(s_{i}^{*},\,\mathbf{s}_{-i}^{B}\bigr).$
@@
A.2 Empirical computation of increasing game complexity
Claim.
Proof.
This is a simple consequence that in games with $k$ different colors of chips, there are $\binom{k}{2}$ possible pairs of chips to transact. Since the distribution of opponents’ valuations is symmetric across chips (except the green numeraire chip), each proposer must consider $O(k^{2})$ different feasible trades. Indeed, if we empirically compute the expected number of myopically rational trades [^15] We define a myopically rational trade to be one where the proposer gains surplus and assigns some probability that at least one opponent would gain surplus by accepting the trade. in the starting configuration of the 2, 3, and 4-chip game, we arrive at 37.1, 120.7, 250.7 respectively. That is the 3-chip game has approximately $\binom{3}{2}$ as many myopically rational trades in expectation as the 2-chip game, and the 4-chip game has approximately $\binom{4}{2}$. This approximation becomes more exact as the number of chips increases due to the reduced relative impact of the numeraire.
A.3 Multi-agent trading algorithm with Bayesian learning

### v4

@@ 6.2.1. Turn-level analysis
Figure 4. Example of player chip value trajectories over nine turns of a trading game (3-chip). To avoid speculation, we do not analyze unobserved strategies. For example, we do not analyze Player 1’s proposal on Turn 4 (an unaccepted trade) or Player 3’s acceptance on Turn 5 (accepting a negative surplus offer).
Example of player chip value trajectories over nine turns of a trading game (3-chip). To avoid speculation, we do not analyze unobserved strategies. For example, we do not analyze Player 1’s proposal on Turn 4 (an unaccepted trade) or Player 3’s acceptance on Turn 5 (accepting a negative surplus offer).
6.2.2. Retrospective analysis
We now evaluate decisions retrospectively, asking whether they were optimal given the opportunities that actually materialized later in the game. At each turn, a player takes one of three actions: proposing, accepting, or declining a trade. We classify each action by its regret properties:
  • No regret: There was no alternative move in a later round that would have yielded higher surplus.
  • Forced regret: A higher-surplus option arose later, but earlier decisions (such as premature trades or overcommitting inventory) prevented the agent from acting on it.
  • Unforced regret: A higher-surplus move was available to take (i.e., no budget or chip count constraints), yet the agent failed to take it.
We analyze how these regret classifications apply to the actions of the different agent types. To avoid speculative modeling of unobserved strategies, we restrict our analysis to a subset of trades where the agent’s intent is clear. First, we only consider myopically rational transactions for proposers, acceptors, or decliners. [^12] That is, only transactions yielding positive surplus. Note that Bayesian agents satisfy this by design. Interpreting irrational trades would require assumptions we cannot credibly test, as they may be driven by factors such as reciprocity, signaling, or cognitive error. Second, when evaluating proposer behavior, we restrict our regret analysis to focus on accepted trades. The act of rejecting an offer could be a mistake, but it could also be strategic signaling. While acceptances can also stem from participant error, they provide observable realized gains that allow for objective regret calculation. Third, “acceptors” include those who intended to accept a trade, not only those who were selected to trade. Recall that this occurs when both other players in a game accept the trade; one is randomly chosen.
Figure 4 shows an example of a player’s valuations as they trade over the course of a single game. On the left, private valuations are shown: Player 3 values blue chips at \$0.30 and red chips at \$0.10. On turn 1, Player 1 proposes to give 9 blue chips to receive 9 red chips in return. Player 3 is an acceptor and ends the turn with 19 blue chips and 1 red chip, netting a \$1.80 surplus. However, in turn 2, Player 2 offers a more lucrative offer: 10 blue chips for 10 red chips, which would give Player 3 a higher surplus of \$2.00 if they could accept it. However, Player 3 does not have enough red chips to clear the Turn 2 trade. Therefore, we characterize their action in Turn 1 as forced regret by the acceptor.
On turn 7, Player 2 makes an unforced regret decline. Player 1 offers to give two green chips for three blue chips. Player 2 has enough chips to clear this trade at this point; if Player 2 were to accept, they would gain $0.5*2-0.3*3=\mathdollar 0.10$. However, they decline the profitable trade and do not recover the profits later in the game. [^13] We do not classify proposals as unforced regret; agents may send intentionally conservative offers to elicit signals on market dynamics.
On turn 8, Player 2 makes a no regret proposal. Player 2 offers to give six red chips to receive eight green chips, resulting in a surplus of $0.5*8-0.6*6=\mathdollar 0.40$. This was the best deal they could have made, as there were no better trade opportunities in the future. Additional details on the regret taxonomy are provided in Appendix D.
Figure 5 counts these actions across agent types and game complexity. Each panel corresponds to a game variant (2, 3, or 4 chips); rows group agents by role (proposer, acceptor, decliner). Green bars indicate no-regret actions, blue bars unforced regret, and red bars forced regret. Because we restrict our analysis to unambiguous actions, these bars reflect the frequency of classifications within the filtered sample. The relative proportions of color within each bar, rather than the absolute bar length, provide the relevant comparison of strategic behavior across agent types.
Figure 5. Counts of actions by regret category across agent types and game complexity. Green is no regret, blue is unforced regret, and red is forced regret. Rows group agents by role (proposer, acceptor, decliner). Note: declining cannot cause forced regret since it does not commit chips.
Counts of actions by regret category across agent types and game complexity. Green is no regret, blue is unforced regret, and red is forced regret. Rows group agents by role (proposer, acceptor, decliner). Note: declining cannot cause forced regret since it does not commit chips.
Several patterns emerge. The proportion of green (no regret) in Bayesian proposals relative to other classifications is high. That is, they consistently identify optimal trades. In contrast, human and LLM proposers show substantial relative unforced regret (blue); more often than their Bayesian counterparts, they leave surplus on the table when better options were available. This is by construction; the Bayesian agents are myopically rational and accept all positive-surplus offers. Figure 2 shows that Bayesian agents achieve the highest realized surplus overall, suggesting that this myopic surplus-seeking policy may be well-suited for maximizing value in this environment.
6.3. Qualitative Survey Feedback
Finally, we examine participants’ subjective experiences through post-game survey responses. Despite increasing game complexity, self-reported cognitive exertion, satisfaction, and performance certainty did not differ significantly across the 2-, 3-, and 4-chip conditions. This effort invariance is consistent with the use of fixed-complexity heuristics (such as the balanced ’equal-volume’ trades observed in Figure 3), which may allow the human players to maintain a stable cognitive load regardless of the underlying state-space complexity.
Participants additionally described their trading strategies in words. Thematic analysis (Braun and Clarke, 2006) of these freeform responses revealed three primary themes: profit maximization (cited in 112 responses), collaborative strategy (56), and color optimization (55). [^14] These distributions are visualized in Figure 9. Responses were double-coded if they addressed multiple themes. Profit maximization was the most frequently cited approach (112 responses), with participants describing goals like maximizing their own chip gains. Collaborative and color-focused strategies appeared at similar rates (56 and 55 responses, respectively). Many participants prioritized fairness and mutual benefit, with explanations such as “I aimed to be fair and reasonable with my trades,” or “I kept my offers low to encourage cautious players to engage.” Others took a more aggressive approach, such as “I traded aggressively early on to secure chips before others could deplete the supply.”
These qualitative accounts align with the observed human tendency toward reciprocal exchanges, where participants request a volume of chips similar to the amount offered. These accounts can also be interpreted alongside the self-reported effort invariance. For example, color-based strategies can help players simplify decision-making and manage their cognitive load. Notably, however, none of these self-reported strategies (collaborative, profit-maximizing, or color-optimizing) correlated with actual game performance.
7. Discussion
We briefly discuss behavioral patterns observed across agent types and offer recommendations for deployment and evaluation.
7.1. Empirical Findings
Human behaviors.
Our study sample comprised online strangers engaging in one-shot, mostly anonymous interactions. There was effectively no reputational risk between players—they did not even know whether they were playing against the same humans in the second game. Given this, a traditional game-theoretic model might predict more extractive behaviors prioritizing individual utility maximization (Rubinstein, 1982; Young, 2007). However, the humans offered trades consistent with fairness norms, i.e., the 1:1 ratio of chips offered to chips requested. After the games, themes like coordination and fairness emerged as primary objectives. Such proposals generated less total surplus than Bayesian agents but were more frequently accepted, though humans still exhibited substantial unforced regret, leaving surplus on the table even when better trades were available. In sum, these behaviors could suggest that humans anticipate that their human counterparts will reject unfair offers, and may be more individually rational than overall surplus would suggest.
These findings reveal a significant tension between traditional game-theoretic predictions and observed human behavior. Even in an environment with anonymity and one-shot interactions, specifically designed to minimize social accountability, human behavior remains conspicuously non-extractive. This also suggests that human objectives are not easily reduced to standard reward functions; rather, strategic reasoning appears deeply rooted in social context and background knowledge that non-human agents may fail to capture. In a mixed-agent future, it is unclear whether these human fairness norms will hold against non-human negotiators who may not share such norms.
LLM behaviors.
Despite the novel domain, some GPT-4o-powered agents achieved human-level surplus with almost no additional prompting. They achieved this surplus through offering generally concessionary proposals —giving more chips than they requested—which were readily accepted by other LLM agents. However, this acceptance-maximizing approach led to high rates of forced regret, as agents committed chips early and foreclosed superior opportunities later in the game.
While the underlying causes of this baseline conservative play style remain opaque, it likely arises from three intersecting factors: (i) instruction-tuning on cooperative, information-sharing dialogues where minimizing friction is implicitly rewarded (Wei et al., 2022); (ii) a tendency toward risk-averse, passive responses exacerbated by a lack of outcome-driven feedback during the task (Ouyang et al., 2022); and (iii) a learned aversion to asymmetric outcomes, where generous offers serve as a heuristic to secure rapid coordination (Fisher et al., 2011).
The implications of this default behavior are significant. In non-cooperative settings such as zero-sum negotiations, this overly concessionary tendency could lead LLMs to systematically over-compromise at baseline, in the absence of explicitly prompted behavioral changes. Unless post-training techniques evolve to explicitly account for strategic context, designers may need to intentionally steer these models to prevent such biases from being exploited by more extractive counterparts in competitive environments.
The performance gap between GPT-4o and Gemini 1.5 Pro may also reflect differences in model development cycles. Gemini 1.5 Pro was released approximately three months prior to GPT-4o, and the rapid pace of capability improvements in this period could account for some of the observed differences, though we cannot isolate model architecture from training data or other factors.
Bayesian behaviors.
Despite their trades being frequently rejected, Bayesian agents consistently captured the highest surplus. Their proposals were predominantly no-regret, reflecting strategic precision that held even as game complexity increased. While impressive, this strong performance (in terms of surplus) required us to write a bespoke statistical algorithm specifically aligned to the game’s incentives. We have no reason to believe these behaviors will effectively translate to real-world negotiations, especially in repeated interactions requiring trust and reciprocity. Extending Bayesian agents to more socially compatible trading would require richer models of preferences, fairness, and adaptive behavior, potentially increasing complexity and decreasing robustness.
7.2. Direct Implications
Beyond the empirical results, a central contribution of this work is demonstrating the limits of outcome-based evaluation for agentic negotiation systems. Historically, models are trained and benchmarked using static, aggregate metrics such as total surplus. Had our analysis stopped at this level, one might reasonably conclude that some LLMs are functional substitutes for human negotiators. However, we find substantial heterogeneity in how these LLMs behave relative to their human counterparts, both in their trading behavior and in the opportunities they later forego. While uncovering these procedural discrepancies in this study required novel evaluation methods that go beyond current, static agentic benchmarks, these procedural differences may be critical in a practitioner’s decision to deploy agentic systems in a real-world scenario.
Additionally, if one were considering deployment in this trading game, it is not clear whether LLMs or Bayesian agents are the better choice. For satisficing performance, we find that LLMs can generalize well to novel negotiation domains with minimal context-specific prompt-engineering. Indeed, additional prompting provided no advantage over the baseline. Both LLMs and Bayesian agents resolve negotiations near-instantaneously and are relatively costless at the margin compared to humans. These efficiency advantages make automated agents appealing for high-volume or time-sensitive negotiations, even when beliefs about the other agents are imperfect.
For optimal performance, tradeoffs are more complex. Human agents appear to trade according to “unwritten” fairness rules. Bayesian agents maximize surplus but incur frequent rejections that could prove costly in repeated or reputation-sensitive settings.
Given the current model capabilities and limitations, an optimal implementation may involve a modular hybrid approach combining human values with the speed and reasoning of automated agents. Here, we can imagine LLMs paired with planning or optimization components, such as Bayesian tools for valuation inference, inventory reasoning, or limited lookahead, to improve strategic foresight. Humans as reviewers can then evaluate the final decisions to ensure they accord with the social context.
Concretely, we recommend that both researchers studying human–AI alignment and practitioners building or deploying AI negotiation agents in collective settings—especially those where agents operate alongside humans or act on their behalf—evaluate (i) outcome metrics, such as deal rate and achieved surplus; (ii) procedural metrics, such as strategic behaviors; and (iii) alignment metrics, assessing whether an agent’s strategy reflects the intent, risk tolerance, and normative expectations of the human counterpart.
7.3. Societal Implications
Our findings also have broader societal implications for the deployment of AI agents in human-facing negotiation and coordination settings. In our study, humans consistently proposed fair trades despite no explicit incentives for doing so. One plausible explanation is that participants understood their counterparts to be other humans who value fairness, coordination, and mutual acceptance, and adjusted their behavior accordingly. In this sense, human negotiation behavior appears to be shaped not only by material incentives but also by beliefs about the counterpart’s identity and normative expectations. This is consistent with much of the behavioral game theory literature (Camerer, 2003).
These dynamics raise important questions for increasingly mixed-agent environments, where such beliefs may no longer be well calibrated. If humans come to recognize that they are negotiating with concessionary LLMs, they may rationally adopt more aggressive strategies. Conversely, repeated interaction with highly extractive, optimization-driven agents may undermine human fairness norms altogether, shifting negotiations toward more adversarial and less cooperative equilibria. In both cases, agents that achieve comparable surplus through fundamentally different negotiation styles may produce unexpected and differential outcomes in new settings.
8. Limitations and Future Work
@@
Our study focuses on a single, stylized negotiation game with static valuations and single-shot interactions, conditions that align well with Bayesian agents’ task-specific optimization and may exaggerate their advantage. While humans and LLMs bring general-purpose reasoning to the task, their broader capabilities in dynamic, multi-round, or reputation-driven negotiations remain unexplored. Furthermore, our LLMs used minimal prompting, reflecting realistic deployment but limiting exploration of more strategic behaviors from alternative prompts or fine-tuning.
@@
Beyond performance metrics, our analysis reveals distinct behavioral profiles but does not disentangle whether these arise from reasoning limitations, social norms, or inductive biases inherent to each system. Future work should investigate how humans perceive and trust AI negotiators with different behaviors. Another promising avenue for future work is to allow agents to communicate with natural language. Finally, developing methods to steer LLM agents towards desired procedural norms represents a significant ongoing challenge and opportunity. To support this agenda, we’ve released our game implementation on Deliberate Lab (Qian et al., 2025b), enabling researchers to benchmark new models against the human, LLM, and Bayesian baselines established here as they emerge.
9. Conclusion
As LLMs begin to match human-level outcomes in social decision-making tasks, the central question has shifted from whether they can perform to how they do so—and whether their strategies are robust, predictable, and normatively acceptable.
This study provides a controlled empirical approach for comparing humans, LLMs, and bespoke statistical agents in a dynamic bargaining setting. We organize our findings around two questions: how these agents perform in terms of achieved surplus, and how they behave in the process of getting there. For RQ1 (performance), Bayesian agents capture the most surplus, while GPT-4o-based agents can achieve human-level surplus. For RQ2 (process), we show that similar aggregate outcomes can arise from sharply different negotiation styles: humans tend toward fairness-constrained exchange, while LLMs default to concessionary, acceptance-seeking proposals.
These procedural differences are likely to matter more as LLM agents move beyond stylized games and static interactions into longer-horizon, higher-stakes interactions embedded in ongoing relationships. Agents that concede too readily invite exploitation, while agents that extract too aggressively risk rejection, mistrust, or norm violations. Because identifying these failure modes depends on strategic analysis rather than outcome comparison, static evaluations can be misleading. Responsible deployment, therefore, requires process-sensitive evaluation that makes strategic posture legible, quantifying missed opportunities and checking alignment with human stakeholders’ goals, risk tolerance, and fairness norms.
10. GenAI Usage Disclosure
LLMs were deployed as one of our populations of study. Portions of Figure 1 were generated using GPT-4o. LLM-assisted tools were used for code editing in data analysis notebooks, proofreading, and light copyediting. Generative AI was not used in other parts of the research process, including literature review, substantial data analysis, or writing. The authors retain full responsibility for all content and conclusions presented herein.
@@ Appendix A Algorithms and Proofs
A.1. Empirical Computation of Increasing Game Complexity
Claim.
Proof.
This is a simple consequence that in games with $k$ different colors of chips, there are $\binom{k}{2}$ possible pairs of chips to transact. Since the distribution of opponents’ valuations is symmetric across chips (except the green numeraire chip), each proposer must consider $O(k^{2})$ different feasible trades. Indeed, if we empirically compute the expected number of myopically rational trades [^15] We define a myopically rational trade to be one where the proposer gains surplus and assigns some probability that at least one opponent would gain surplus by accepting the trade. in the starting configuration of the 2, 3, and 4-chip game, we arrive at 37.1, 120.7, 250.7 respectively. That is, the 3-chip game has approximately $\binom{3}{2}$ as many myopically rational trades in expectation as the 2-chip game, and the 4-chip game has approximately $\binom{4}{2}$ times as many. This approximation becomes more precise as the number of chips increases, since the numeraire’s relative impact is reduced.
A.2. Proof of No Dominant Strategy
@@
Construct two different scenarios of other agents’ strategies and valuations, denoted by $\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$. These scenarios can differ in (i) how other agents value the goods $\{v_{jg}:j\neq i\}$, and (ii) the demands or offers they make (i.e. how $\mathbf{s}_{-i}$ translates into allocations $a_{jg}$).
@@
  • In scenario $B$, suppose the others behave differently (e.g., they no longer place much value on $g^{*}$ but highly value some other good). Now, a different specialized strategy $s_{i}^{B}$ might yield a strictly higher payoff (because the trading strategy for $g^{*}$ in scenario $A$ is no longer beneficial here).
Contradiction of dominance. Since $A$ and $B$ are feasible opponent profiles, no single $s_{i}^{*}$ can weakly dominate both tailored responses $s_{i}^{A}$ and $s_{i}^{B}$ across them. In particular,
(1) $\exists\,\mathbf{s}_{-i}^{A}\;:\;$ $u_{i}\!\left(s_{i}^{A},\mathbf{s}_{-i}^{A}\right)\;>\;u_{i}\!\left(s_{i}^{*},\mathbf{s}_{-i}^{A}\right),$
(2) $\exists\,\mathbf{s}_{-i}^{B}\;:\;$ $u_{i}\!\left(s_{i}^{B},\mathbf{s}_{-i}^{B}\right)\;>\;u_{i}\!\left(s_{i}^{*},\mathbf{s}_{-i}^{B}\right).$
@@
A.3. Multi-Agent Trading Algorithm with Bayesian Learning